Suppose that the birth weight of babies X has a normal distribution with mean µ = 2.4 kg, and standard deviation σ =0.5 kg. A baby chosen at random. Find The birth weight C if PX>x=0.60.
1
Expert's answer
2019-11-11T13:20:09-0500

"X\sim N(\mu, \sigma^2)"

Then


"Z={X-\mu \over \sigma}\sim N(0,1)"

Given that "\mu=2.4\ kg, \sigma=0.5\ kg"

"P(X>x)=0.60"


"P(X>x)=1-P(X\leq x)=1-P(Z\leq {x-2.4 \over 0.5})=0.60"

"P(Z\leq {x-2.4 \over 0.5})=1-0.60=0.40"

"{x-2.4 \over 0.5}\approx-0.253348"

"x\approx0.5(-0.253348)+2.4\approx2.2733"

"\text{Birth weight}=2.2733\ kg"